OpenStudy (diana):

Find the derivative of y with respect to the appropriate variable. y = s(1-s^2)^(1/2) + cos^-1(s)

OpenStudy (diana):

\[y = s \sqrt{1-s ^{2}} + \cos^{-1} s\]

OpenStudy (anonymous):

Split into 2 integrals In the first one, substitute u = 1 - s^2 Check integral table for the second one!